A line passes through the points (5, 3) and (2, 5). a) Find the equation of the line b) Graph the line using the above two points. please show how to s

Algebra ->  Finance -> SOLUTION: Question 4: A line passes through the points (5, 3) and (2, 5). a) Find the equation of the line b) Graph the line using the above two points. please show how to s      Log On


   



Question 1015596: Question 4: A line passes through the points (5, 3) and (2, 5).
a) Find the equation of the line
b) Graph the line using the above two points.

please show how to solve

Answer by Alan3354(69443)

Question 4: A line passes through the points (5, 3) and (2, 5).
a) Find the equation of the line
slope m = diffy/diffx
m = 2/-3 = -2/3
Use y-y1 = m*(x-x1) where (x1,y1) is either point.
y-3 = (-2/3)*(x-5)
======================
b) Graph the line using the above two points.
Plot the points, draw a line thru them.